A person has a mass of 65 kg. The coefficient of static friction µs, between their shoes and the floor = 0.8. Calculate the force required to slide them across the floor. 


1
Expert's answer
2020-11-20T09:30:21-0500

The magnitude of friction force is defined as:


"F_{fr} = \mu N"

where "\mu = 0.8" is the coefficient of static friction and "N" is the magnitude of the normal force.

Since the person does not move in the vertical direction, then


"N = mg"


Here "m = 65kg" is the person's mass and "g = 9.81m\/s^2" is the gravitational acceleration.

Then, obtain:


"F_{fr} = \mu mg\\
F_{fr} = 0.8\cdot 65\cdot 9.81 = 510.12 \space N"

In order to slide them accros the floor, one should apply the force greater or equal to "F_{fr}".


Answer. "F \ge 510.12\space N".